# Variable Inertia Model Predictive Control for Fast Bipedal Maneuvers

### v1

@@ IV-B Simulation Results
To investigate the effect of centroidal rotational inertia in MPC, we compared three scenarios with different inertia update rules: (1) employing constant inertia at a nominal configuration over the MPC horizon, (2) using constant inertia over the MPC horizon but updating it at every MPC iteration, and (3) applying the CCINN to predict the varying inertia over the MPC horizon (proposed method), as depicted in Fig. 3. Thanks to the convex formulation and the efficacy of CCINN, the average MPC solve time was approximately $1.4969$ $\mathrm{m}\mathrm{s}$, as illustrated in Fig. 4. For cases (1) and (2), we initialized the MPC with the torso’s orientation and velocity at every MPC update, similar to the baseline controller. For case (3), we initialized the MPC with the WBO and WBO velocity.
@@
Figure 6: Torso Orientation vs. WBO. Comparison of torso orientation (blue line) and WBO (black line) in the circular walking scenario (0.9 $\mathrm{m}\mathrm{/}\mathrm{s}$ and 0.9 $\mathrm{r}\mathrm{a}\mathrm{d}\mathrm{/}\mathrm{s}$).
V DISCUSSION AND CONCLUSIONS
In this work, we have introduced a MPC framework for agile and robust bipedal locomotion. By leveraging the variable centroidal inertia model and WBO, we developed a VI-MPC framework that generates dynamically consistent contact wrenches. We have demonstrated the efficacy of our approach across various locomotion scenarios, where the robot achieved increased step lengths, walking speeds, and turning speeds. In future work, we will explore how this framework can be extended to various humanoid loco-manipulation scenarios.
@@ ACKNOWLEDGMENT
This work was supported by ONR Grant# N000142212204.

### v2

@@ IV-B Simulation Results
To investigate the effect of centroidal rotational inertia in MPC, we compared our proposed approach against two baseline approaches with different inertia update rules: (1) using CCINN to predict the varying inertia over the MPC horizon (proposed approach), (2) employing constant inertia at a nominal configuration throughout the MPC horizon (baseline 1), and (3) using constant inertia over the MPC horizon but updating it at each MPC iteration (baseline 2), as depicted in Fig. 3. Thanks to the convex formulation and the efficacy of CCINN, the average MPC solve time was approximately $1.4969$ $\mathrm{m}\mathrm{s}$, as shown in Fig. 4. For the baseline approaches, we initialized the MPC with the torso’s orientation and velocity at every MPC update. For our proposed approach, we initialized the MPC with the WBO and WBO velocity.
@@
Figure 6: Torso Orientation vs. WBO. Comparison of torso orientation and WBO in the circular walking scenario (0.9 $\mathrm{m}\mathrm{/}\mathrm{s}$ and 0.9 $\mathrm{r}\mathrm{a}\mathrm{d}\mathrm{/}\mathrm{s}$). Red arrows represent the initial heading, and the blue arrows indicate the current heading with velocity commands.
V CONCLUSIONS
In this work, we introduced a MPC framework for agile and robust bipedal locomotion. By leveraging the variable centroidal inertia model and WBO, we developed a VI-MPC framework that generates dynamically consistent contact wrenches. We have demonstrated the efficacy of our approach across various locomotion scenarios, where the robot achieved increased step lengths, walking speeds, and turning speeds. Future work will focus on extensive hardware experiments with DRACO $3$ to further validate the proposed method. Additionally, we will explore the application of this framework to diverse humanoid loco-manipulation tasks that involve substantial changes in inertia.
@@ ACKNOWLEDGMENT
This work was supported by the Office of Naval Research (ONR), Award No. N00014-22-1-2204.